OpenStudy (anonymous):

Factor Completely: x2 − 36

OpenStudy (anonymous):

(x+6)(x-6)

hero (hero):

(x+6)(x-6)

OpenStudy (lalaly):

difference between two squares x^2-a^2 = (x-a)(x+a) x^2-36= (x+6)(x-6)

OpenStudy (anonymous):

x^2-36=0 (x+6)(x-6)=0 x can equal -6 or 6.

OpenStudy (anonymous):

\[x^2 - y^2=(x+y)(x-y)\]

OpenStudy (anonymous):

The diff. between Two squares is quite simple. x^2 - 6^2. Then it is (x-6)(x+6)